OpenStudy (anonymous):

Exponential Exponent Expression Thanks (3b^3)^4

OpenStudy (anonymous):

(3b)^12

OpenStudy (anonymous):

@Allenismyname did u get it?!

OpenStudy (anonymous):

I don't think that is right.

OpenStudy (anonymous):

I got 81b^12

Nnesha (nnesha):

\[\huge\rm (ab)^m =a^m b^m\] a b both r raising to the m power

Nnesha (nnesha):

i just want to ask one question how did you get 81 ? :=)

OpenStudy (anonymous):

3^4=81

Nnesha (nnesha):

yes right good job!